# Decoding Open-Ended Information Seeking Goals from Eye Movements in Reading

### v2

@@ E.1 Question Selection
Model
New Item
Same Span
New Participant
Same Span
New Item & Participant
Same Span
Chance / Majority Baseline $50.1\pm 0.4_{\text{+++}}$ $50.0\pm 0.6_{\text{+++}}$ $50.7\pm 2.0_{\text{++}}$
Question Similarity to RT-Weighted Passage $49.3\pm 0.6_{\text{+++}}$ $50.6\pm 0.8_{\text{+++}}$ $49.9\pm 2.2_{\text{++}}$
RT Similarity to Question-Word Similarities $50.9\pm 0.6_{\text{+++}}$ $51.0\pm 0.6_{\text{+++}}$ $53.0\pm 1.5_{\text{+}}$
Haller RNN (Haller et al., 2022) $53.2\pm 0.4^{**}_{\text{+++}}$ $53.4\pm 0.7^{**}_{\text{+++}}$ $52.2\pm 1.7_{\text{+}}$
RoBERTEye-Fixations (Shubi et al., 2024) $\textbf{57.3}\pm\textbf{0.9}^{***}$ $\textbf{59.7}\pm\textbf{0.5}^{***}$ $\textbf{58.6}\pm\textbf{1.9}^{**}$
DalEye-LLaVA $51.0\pm 0.3_{\text{+++}}$ $50.7\pm 0.9_{\text{+++}}$ $49.8\pm 1.5_{\text{+++}}$
DalEye-Llama $50.9\pm 0.4_{\text{+++}}$ $52.5\pm 0.6^{*}_{\text{+++}}$ $53.2\pm 1.2_{\text{+}}$
E.1.1 DalEye-Llama Input Variations
@@
In Table˜11 we present the results of using either one of the two proposed task prompts and three input representations - fixation-level, word-level, and a combined fixation- and word-level representation (see Section˜B.2.1 for the full inputs). Across all three task variations, Task prompt #1 with fixation-level input yields the best accuracy. Both the alternative task prompt and the word-level/combined inputs underperform this setting, where the choice of representation matters more than the task prompt.
@@
Task Prompt Representation All Spans
Diff Span
Same Span
@@
Task prompt #1 Fixation-level $\textbf{39.2}\pm\textbf{0.9}^{***}$ $\textbf{61.4}\pm\textbf{0.9}^{***}$ $\textbf{52.5}\pm\textbf{0.4}^{***}$
@@
Task prompt #2 Fixation-level $37.2\pm 0.6^{***}_{\text{+++}}$ $60.3\pm 0.6^{***}$ $51.0\pm 0.3_{\text{+}}$
@@
E.2 Question Reconstruction
Example generations for DalEye-Llama are presented in Figure˜7. Question reconstruction evaluations of the DalEye-Llama model for (1) the identity of the generated question word, (2) the semantic category of the question (Li & Roth, 2002), (3) similarity using BERTScore, and (4) downstream QA accuracy based on the answer selection of a multiple-choice question answering model. DalEye-Llama performance is benchmarked against two types of human composed questions (for a different critical span and for the same span) and against arbitrary questions generated with GPT-4o and Llama 3.1 in Table˜12 (same as Figure˜4, in numerical format). Presented are means with 95% confidence intervals (bootstrapped, $n=1000$) for three generalization levels to new participants and new readers. The highest score in each combination of evaluation and generalization level is marked in bold. In Table˜13 we present the results of DalEye-Llama compared to the word-level and combined word- and fixation-level input representations, and for the task prompt variation.
True Question New Item Generated Questions New Participant Generated Questions What is among the primary attractions of Lemnos? (A) What is one thing to note about the eponymous village? (A) What is among the primary attractions of Lemnos? (B) Who is quoted as saying that hotels benefit from the increase in the number of tourists in the area? (B) What will happen in spring 2013? (C) What does the quote convey about the current economic situation in Greece? (C) Why does Lemnos have to pay the first round of tax increases? (D) What is true of the other Greek islands? (D) What is one reason why the wealthy tend to spend more time on the islands?
Figure 7: An example showing a paragraph, the corresponding ground truth question, and four DalEye-Llama generated questions from the New Item Regime and four from the New Participant Regime. Each generated question reflects a different outcome from the QA model, as indicated by the answer choice (A–D) selected by the model. The possible answers, structured according to the STARC annotation guidelines (Berzak et al., 2020), to the ground truth question were: A) (correct answer) A large number of cultural destinations. B) (miscomprehension of the critical span) Beaches constantly full of locals and tourists. C) (incorrect, related to a different span) 3,000 luxury hotels. D) (no support in the passage) Highly-regarded restaurants.
Table 12: Test evaluations of the DalEye-Llama model benchmarked against two types of human composed questions (for a different critical span and for the same span) and against arbitrary questions generated with GPT-4o and Llama 3.1.
New Participant New Item
New Participant & Item
Question word (Kappa)
@@ E.2 Question Reconstruction
Incorrect human question - same critical span $087\pm 0.014 0.0$$068\pm 0.013$ $\mathbf{0.107\pm 0.040}$
GPT-4o arbitrary question $020\pm 0.008 0.0$$25\pm 0.007 0.0$$016\pm 0.019$
Llama 3.1 arbitrary question $346\pm 0.011 0.0$$12\pm 0.005 0.0$$003\pm 0.013$
DalEye-Llama question $\mathbf{0.478\pm 0.011}$ $069\pm 0.008 0.0$$052\pm 0.022$
Question category (Kappa)
Incorrect human question - different critical span $0.097\pm 0.009$ $\mathbf{0.107\pm 0.009}$ $\mathbf{0.133\pm 0.028}$
Incorrect human question - same critical span $063\pm 0.014 0.0$$046\pm 0.013$ $0.104\pm 0.040$
GPT-4o arbitrary question $005\pm 0.006 0.0$$18\pm 0.006 0.0$$022\pm 0.019$
Llama 3.1 arbitrary question $250\pm 0.011 0.0$$28\pm 0.009 0.0$$024\pm 0.026$
DalEye-Llama question $\mathbf{0.299\pm 0.011}$ $074\pm 0.010 0.0$$045\pm 0.028$
@@
Incorrect human question - different critical span $0.291\pm 0.002$ $0.291\pm 0.002$ $0.286\pm 0.007$
Incorrect human question - same critical span $0.384\pm 0.004$ $\mathbf{0.380\pm 0.004}$ $\mathbf{0.379\pm 0.011}$
GPT-4o arbitrary question $0.212\pm 0.002$ $0.217\pm 0.002$ $0.210\pm 0.006$
Llama 3.1 arbitrary question $0.523\pm 0.005$ $0.299\pm 0.003$ $0.309\pm 0.007$
DalEye-Llama question $\mathbf{0.602\pm 0.006}$ $0.315\pm 0.003$ $0.314\pm 0.007$
@@
Incorrect human question - different critical span $0.604\pm 0.001$ $0.603\pm 0.001$ $0.601\pm 0.003$
Incorrect human question - same critical span $0.651\pm 0.002$ $\mathbf{0.650\pm 0.002}$ $\mathbf{0.654\pm 0.006}$
GPT-4o arbitrary question $0.413\pm 0.005$ $0.398\pm 0.005$ $0.313\pm 0.013$
Llama 3.1 arbitrary question $0.723\pm 0.003$ $0.617\pm 0.001$ $0.617\pm 0.003$
DalEye-Llama question $\mathbf{0.774\pm 0.003}$ $0.631\pm 0.002$ $0.628\pm 0.004$

### v3

@@ G.1.1 Test and Validation Results across Subtasks and Evaluation Regimes
Model New Participant Same Span New Text Same Span New Text & Participant Same Span
Chance / Majority Baseline $50.0\pm 0.6_{\text{+++}}$ $50.1\pm 0.4_{\text{+++}}$ $50.7\pm 2.0_{\text{++}}$
Question Similarity to RT-Weighted Passage $50.6\pm 0.8_{\text{+++}}$ $49.3\pm 0.6_{\text{+++}}$ $49.9\pm 2.2_{\text{++}}$
RT Similarity to Question-Word Similarities $51.0\pm 0.6_{\text{+++}}$ $50.9\pm 0.6_{\text{+++}}$ $53.0\pm 1.5_{\text{+}}$
Haller RNN (Haller et al., 2022) $53.4\pm 0.7^{**}_{\text{+++}}$ $53.2\pm 0.4^{**}_{\text{+++}}$ $52.2\pm 1.7_{\text{+}}$
RoBERTEye-Fixations (Shubi et al., 2024) $\textbf{59.7}\pm\textbf{0.5}^{***}$ $\textbf{57.3}\pm\textbf{0.9}^{***}$ $\textbf{58.6}\pm\textbf{1.9}^{**}$
DalEye-LLaVA $50.7\pm 0.9_{\text{+++}}$ $51.0\pm 0.3_{\text{+++}}$ $49.8\pm 1.5_{\text{+++}}$
DalEye-Llama $52.5\pm 0.6^{*}_{\text{+++}}$ $50.9\pm 0.4_{\text{+++}}$ $53.2\pm 1.2_{\text{+}}$
G.1.2 Evaluation of Generative Models on Goal Selection
For the generative models, goal selection is performed by computing the model-assigned log-likelihood of each candidate question given the input and selecting the question with the highest log-likelihood. As GPT-4o-mini and Gemini 3 do not allow extracting input token logits, we do not evaluate them on the question selection task.
Table 12: Goal selection test accuracy aggregated over 10 cross-validation splits, with 95% confidence intervals. We test for differences between models using a linear mixed effects model with random intercepts and slopes for participants and paragraphs: $is\_correct\sim model+(model\mid participant)+(model\mid paragraph)$. Model accuracy is compared to the chance performance of the discriminative models. Significant gains over this baseline are marked with ‘**’ $p<0.01$, and ‘***’ $p<0.001$. The best performing model in each evaluation regime is marked in bold.
Model Type Model All Different Spans Same Span
Discriminative Chance / Majority Baseline 33.0 $\pm$ 0.4 55.3 $\pm$ 0.4 49.9 $\pm$ 0.4
Generative DalEye-LLaVA 33.5 $\pm$ 0.3 57.0 $\pm$ 0.7∗∗ 49.4 $\pm$ 0.4
[…]
The results are presented in Table˜12. DalEye-Llama outperforms DalEye-LLaVA, where the latter is at chance level except for the Different Spans condition. The advantage of the discriminative models over the generative models is not surprising, as differently from the discriminative models, the generative models are not explicitly trained on the question selection task.
G.1.3 RoBERTEye-Fixations Feature Ablation
To quantify the contribution of different eye-movement and textual features to the best performing discriminative model, RoBERTEye-Fixations, we conducted a feature ablation study. Here, we repeat the main experiment using the same training procedure and hyperparameter search space, ablating each of three feature groups: fixation-level eye movement features (e.g. current fixation duration, next saccade direction), word-level eye movement features (reading measures aggregated for each word, e.g. sum of all fixation durations on the word), and linguistic features of the words (e.g. word frequency, surprisal). This allows us to evaluate the contributions of two different eye movement representations (fixation-level and word-level) and the importance of enabling interactions of eye movement features with linguistic word characteristics. We further divide the fixation-level features into two groups: fixations and saccades, and evaluate the contribution of each group. Furthermore, we also ablate the total paragraph reading time feature. To complement this analysis, we also evaluate two additional representations: one with only linguistic features and another that includes linguistic features and total paragraph reading time feature, a measure that doesn’t require an eyetracker. Table˜3 presents the features of each feature group.
The feature ablation analysis results are reported in Table˜13. We find that removing any of the feature groups reduces performance numerically across all evaluation settings except for ablating the text features or the saccade features in the Same Span regime. In the Different Spans evaluation regime, ablating any feature group, except for the word-level eye movement features, results in a statistically significant drop in performance. Notably, the drop in performance when removing either total paragraph reading time or text features indicates that the model exploits information beyond eye movement features to differentiate between questions tied to different critical spans. By contrast, in the Same Span evaluation regime, as candidate questions refer to the same critical span, finer-grained information than global reading times or text properties, such as fixation durations, is necessary. The largest and only statistically significant drop in performance is observed when both fixation-level feature groups are ablated jointly, with most of the effect driven by the non-saccade fixation features. Overall, the relatively small effect sizes across ablations and the similar performance of the ablations to the linguistic-features-only and linguistic features plus paragraph reading time representations suggest redundancy among feature groups, and that a substantial portion of the signal is already possibly captured by the fixation sequence via the alignment of fixations to word embeddings.
Table 13: Feature Ablation Results for RoBERTEye-Fixations. The best performing model in each evaluation regime is marked in bold. Significant drops compared to the best model are marked with ‘+‘ ($p<0.05$) and ‘++‘ ($p<0.01$) in subscript.
All Different Spans Same Span
RoBERTEye-Fixations $\mathbf{49.2}$$\pm$ $\mathbf{0.4}$ $\mathbf{71.0}$$\pm$ $\mathbf{0.4}$ 57.1 $\pm$ 0.6
   – Fixation-Level Features 48.2 $\pm$ 0.5+ 70.0 $\pm$ 0.5+ 55.9 $\pm$ 0.8+
    – Fixation Features 48.4 $\pm$ 0.5 70.1 $\pm$ 0.6+ 56.4 $\pm$ 0.6
    – Saccade Features 48.6 $\pm$ 0.6 69.8 $\pm$ 0.5++ $\mathbf{57.2}$$\pm$ $\mathbf{0.7}$
   – Word-Level Features 48.7 $\pm$ 0.6 70.3 $\pm$ 0.4 56.7 $\pm$ 0.7
   – Linguistic Text Features 48.6 $\pm$ 0.6 69.8 $\pm$ 0.5++ $\mathbf{57.2}$$\pm$ $\mathbf{0.7}$
   – Paragraph Reading Time 48.4 $\pm$ 0.5 69.5 $\pm$ 0.4++ 56.9 $\pm$ 0.7
   Only Linguistic Text Features 48.6 $\pm$ 0.6 69.9 $\pm$ 0.4+ 56.6 $\pm$ 0.7
    Only Linguistic Text Features + Paragraph Reading Time 48.2 $\pm$ 0.5+ 69.7 $\pm$ 0.5++ 56.2 $\pm$ 0.7
G.2 Goal Reconstruction
G.2.1 Numerical Results with Statistical Tests
Tables˜14 and 15 present the goal reconstruction evaluations of Figure˜5 in numerical format. Presented are question reconstruction evaluations of DalEye-Llama, DalEye-GPT and Gemini (zero-shot and few-shot) for (1) the identity of the generated question word, (2) UIUC semantic category of the question, (3) BLEU score, (4) BERTScore, (5) downstream QA accuracy based on the answer selection of a multiple-choice QA model. We additionally provide evaluations for DalEye-LLaVA. The models are benchmarked against six baselines that do not include eye movements. Two baselines are human-composed questions for a different and the same critical span. Four additional baselines are LLM-generated questions: an arbitrary question about the text generated with Gemini 3, and questions from text-only variants of DalEye-Llama, DalEye-GPT and DalEye-LLaVA, finetuned on the question decoding task using only the text.
Table 14: Goal reconstruction evaluations on the Question Word accuracy, Question Category accuracy and BLEU measures. Presented are intercept coefficients with 95% confidence intervals from a linear mixed effects model with random intercepts for participants and paragraphs: $measure\sim 1+(1\mid participant)+(1\mid paragraph)$ for eye movement models and the $measure\sim 1+(1\mid paragraph)$ for the baselines. The highest score in each combination of evaluation measure and evaluation regime is marked in bold. Model comparisons were conducted using ${measure}\sim{model}+(1\mid participant)+(1\mid paragraph)$. Significant drops compared to the best model are marked in superscript with ’*’ $p<0.05$, ’**’ $p<0.01$, and ’***’ $p<0.001$. Significant improvements over corresponding baselines are marked in subscript with ’+’ $p<0.05$, ’++’ $p<0.01$, and ’+++’ $p<0.001$.
New Participant New Item New Participant & Item
Question Word accuracy
Incorrect Human (different critical span) $0.416\pm 0.022^{***}$ $0.407\pm 0.022^{***}$ $0.403\pm 0.024^{***}$
Incorrect Human (same critical span) $0.370\pm 0.019^{***}$ $0.370\pm 0.019^{***}$ $0.401\pm 0.021^{***}$
Text-only Llama 3.1 $0.591\pm 0.014^{***}$ $0.540\pm 0.018$ $0.540\pm 0.020$
DalEye-Llama $0.674\pm 0.013^{***}_{\text{+++}}$ $0.545\pm 0.017$ $0.548\pm 0.023$
Text-only GPT-4o-mini $0.561\pm 0.014^{***}$ $0.450\pm 0.014^{***}$ $0.479\pm 0.020^{***}$
DalEye-GPT $0.617\pm 0.014^{***}_{\text{+++}}$ $0.495\pm 0.015^{***}_{\text{+++}}$ $0.496\pm 0.023^{***}$
Arbitrary Gemini 3 $0.318\pm 0.013^{***}$ $0.327\pm 0.013^{***}$ $0.328\pm 0.019^{***}$
Gemini zero-shot $0.382\pm 0.012^{***}_{\text{+++}}$ $0.378\pm 0.013^{***}_{\text{+++}}$ $0.393\pm 0.022^{***}_{\text{+++}}$
Gemini few-shot $\mathbf{0.713\pm 0.012}_{\text{+++}}$ $0.395\pm 0.013^{***}_{\text{+++}}$ $0.410\pm 0.022^{***}_{\text{+++}}$
Text-only LLaVA OneVision $0.538\pm 0.017^{***}$ $0.499\pm 0.019^{***}$ $0.498\pm 0.021^{***}$
DalEye-LLaVA $0.558\pm 0.018^{***}_{\text{+++}}$ $\mathbf{0.548\pm 0.019}_{\text{+++}}$ $\mathbf{0.548\pm 0.024}_{\text{+++}}$
Question Category accuracy
Incorrect Human (different critical span) $0.410\pm 0.021^{***}$ $0.410\pm 0.021^{***}$ $0.423\pm 0.024^{*}$
Incorrect Human (same critical span) $0.349\pm 0.019^{***}$ $0.349\pm 0.019^{***}$ $0.357\pm 0.021^{***}$
Text-only Llama 3.1 $0.535\pm 0.014^{***}$ $0.454\pm 0.017^{***}$ $0.453\pm 0.021$
DalEye-Llama $\mathbf{0.558\pm 0.014}_{\text{+++}}$ $0.450\pm 0.017^{***}$ $0.428\pm 0.023^{***}$
Text-only GPT-4o-mini $0.458\pm 0.013^{***}$ $0.430\pm 0.014^{***}$ $0.429\pm 0.020^{**}$
DalEye-GPT $0.550\pm 0.013_{\text{+++}}$ $\mathbf{0.488\pm 0.015}_{\text{+++}}$ $0.454\pm 0.022_{\text{++}}$
Arbitrary Gemini 3 $0.345\pm 0.013^{***}$ $0.367\pm 0.014^{***}$ $0.355\pm 0.019^{***}$
Gemini zero-shot $0.327\pm 0.012^{***}$ $0.335\pm 0.012^{***}$ $0.323\pm 0.021^{***}$
Gemini few-shot $0.526\pm 0.014^{***}_{\text{+++}}$ $0.418\pm 0.013^{***}_{\text{+++}}$ $0.439\pm 0.022^{**}_{\text{+++}}$
Text-only LLaVA OneVision $0.408\pm 0.013^{***}$ $0.424\pm 0.018^{***}$ $0.412\pm 0.021^{***}$
DalEye-LLaVA $0.445\pm 0.015^{***}_{\text{+++}}$ $0.466\pm 0.019^{***}_{\text{+++}}$ $\mathbf{0.460\pm 0.024}_{\text{+++}}$
BLEU
Incorrect Human (different critical span) $0.219\pm 0.006^{***}$ $0.219\pm 0.006^{***}$ $0.219\pm 0.007^{***}$
Incorrect Human (same critical span) $0.298\pm 0.009^{***}$ $\mathbf{0.298\pm 0.009}$ $\mathbf{0.300\pm 0.009}$
Text-only Llama 3.1 $0.454\pm 0.006^{***}$ $0.236\pm 0.004^{***}$ $0.238\pm 0.005^{***}$
DalEye-Llama $0.556\pm 0.008^{***}_{\text{+++}}$ $0.253\pm 0.005^{***}_{\text{+++}}$ $0.244\pm 0.006^{***}$
Text-only GPT-4o-mini $0.452\pm 0.006^{***}$ $0.240\pm 0.004^{***}$ $0.237\pm 0.006^{***}$
DalEye-GPT $0.514\pm 0.008^{***}_{\text{+++}}$ $0.257\pm 0.004^{***}_{\text{+++}}$ $0.256\pm 0.007^{***}_{\text{+++}}$
Arbitrary Gemini 3 $0.191\pm 0.004^{***}$ $0.194\pm 0.005^{***}$ $0.184\pm 0.006^{***}$
Gemini zero-shot $0.246\pm 0.005^{***}_{\text{+++}}$ $0.247\pm 0.005^{***}_{\text{+++}}$ $0.242\pm 0.009^{***}_{\text{+++}}$
Gemini few-shot $\mathbf{0.644\pm 0.009}_{\text{+++}}$ $0.272\pm 0.005^{***}_{\text{+++}}$ $0.268\pm 0.008^{***}_{\text{+++}}$
Text-only LLaVA OneVision $0.304\pm 0.005^{***}$ $0.210\pm 0.004^{***}$ $0.207\pm 0.005^{***}$
DalEye-LLaVA $0.265\pm 0.005^{***}$ $0.231\pm 0.006^{***}_{\text{+++}}$ $0.231\pm 0.007^{***}_{\text{+++}}$
Table 15: Goal reconstruction evaluations on the BERTScore and QA accuracy measures. Presented are intercept coefficients with 95% confidence intervals from a linear mixed effects model with random intercepts for participants and paragraphs: $measure\sim 1+(1\mid participant)+(1\mid paragraph)$ for eye movement models and the $measure\sim 1+(1\mid paragraph)$ for the baselines. The highest score in each combination of evaluation measure and evaluation regime is marked in bold. Model comparisons were conducted using ${measure}\sim{model}+(1\mid participant)+(1\mid paragraph)$. Significant drops compared to the best model are marked in superscript with ’*’ $p<0.05$, ’**’ $p<0.01$, and ’***’ $p<0.001$. Significant improvements over corresponding baselines are marked in subscript with ’+’ $p<0.05$, ’++’ $p<0.01$, and ’+++’ $p<0.001$.
New Participant New Item New Participant & Item
BERTScore
Incorrect Human (different critical span) $0.604\pm 0.004^{***}$ $0.603\pm 0.004^{***}$ $0.602\pm 0.004^{***}$
Incorrect Human (same critical span) $0.651\pm 0.004^{***}$ $\mathbf{0.651\pm 0.004}$ $\mathbf{0.651\pm 0.004}$
Text-only Llama 3.1 $0.723\pm 0.003^{***}$ $0.617\pm 0.003^{***}$ $0.618\pm 0.003^{***}$
DalEye-Llama $0.776\pm 0.004^{***}_{\text{+++}}$ $0.631\pm 0.003^{***}_{\text{+++}}$ $0.626\pm 0.004^{***}_{\text{+++}}$
Text-only GPT-4o-mini $0.722\pm 0.003^{***}$ $0.619\pm 0.003^{***}$ $0.618\pm 0.004^{***}$
DalEye-GPT $0.755\pm 0.004^{***}_{\text{+++}}$ $0.630\pm 0.003^{***}_{\text{+++}}$ $0.627\pm 0.004^{***}_{\text{++}}$
Arbitrary Gemini 3 $0.611\pm 0.003^{***}$ $0.612\pm 0.003^{***}$ $0.605\pm 0.003^{***}$
Gemini zero-shot $0.628\pm 0.003^{***}_{\text{+++}}$ $0.629\pm 0.003^{***}_{\text{+++}}$ $0.627\pm 0.005^{***}_{\text{+++}}$
Gemini few-shot $\mathbf{0.818\pm 0.005}_{\text{+++}}$ $0.642\pm 0.003^{***}_{\text{+++}}$ $0.644\pm 0.005^{***}_{\text{+++}}$
Text-only LLaVA OneVision $0.647\pm 0.003^{***}$ $0.595\pm 0.003^{***}$ $0.591\pm 0.003^{***}$
DalEye-LLaVA $0.638\pm 0.003^{***}$ $0.618\pm 0.003^{***}_{\text{+++}}$ $0.617\pm 0.004^{***}_{\text{+++}}$
QA accuracy
Incorrect Human (different critical span) $0.495\pm 0.018^{***}$ $0.490\pm 0.017^{***}$ $0.490\pm 0.020^{***}$
Incorrect Human (same critical span) $0.677\pm 0.021^{***}$ $0.677\pm 0.021$ $0.686\pm 0.022$
Text-only Llama 3.1 $0.684\pm 0.012^{***}$ $0.609\pm 0.014^{***}$ $0.616\pm 0.020^{***}$
DalEye-Llama $0.763\pm 0.011^{***}_{\text{+++}}$ $0.648\pm 0.014^{***}_{\text{+++}}$ $0.644\pm 0.023^{**}_{\text{+}}$
Text-only GPT-4o-mini $0.679\pm 0.012^{***}$ $0.619\pm 0.013^{***}$ $0.608\pm 0.020^{***}$
DalEye-GPT $0.738\pm 0.011^{***}_{\text{+++}}$ $0.658\pm 0.013^{***}_{\text{+++}}$ $0.644\pm 0.023^{***}$
Arbitrary Gemini 3 $0.635\pm 0.015^{***}$ $0.636\pm 0.015^{***}$ $0.614\pm 0.020^{***}$
Gemini zero-shot $0.674\pm 0.013^{***}_{\text{+++}}$ $0.664\pm 0.013^{***}_{\text{+++}}$ $\mathbf{0.690\pm 0.022}_{\text{+++}}$
Gemini few-shot $\mathbf{0.804\pm 0.010}_{\text{+++}}$ $\mathbf{0.683\pm 0.012}_{\text{+++}}$ $0.671\pm 0.023^{*}_{\text{+++}}$
Text-only LLaVA OneVision $0.632\pm 0.012^{***}$ $0.636\pm 0.014^{***}$ $0.612\pm 0.019^{***}$
DalEye-LLaVA $0.623\pm 0.013^{***}$ $0.610\pm 0.016^{***}$ $0.579\pm 0.024^{***}$
G.2.2 Example Generations
Example generations for DalEye-Llama, DalEye-GPT, Gemini zero-shot, and Gemini few-shot are presented in Table˜16. The complete set of model outputs for all three evaluation regimes, including the baselines, together with the corresponding annotations, is released in the project’s GitHub repository under generated_questions.csv.
Table 16: An example showing a paragraph, the corresponding ground truth question, and, for DalEye-Llama, Gemini (zero-shot and few-shot), and DalEye-GPT, four generated questions from the New Text Regime and four from the New Participant Regime. Each generated question reflects a different outcome from the QA model, as indicated by the answer choice (A–D) selected by the model. "(no example)" indicates that the model did not generate a question for the corresponding answer choice. The possible answers, structured according to the STARC annotation guidelines (Berzak et al., 2020), to the ground truth question were: A) (correct answer) A large number of cultural destinations. B) (miscomprehension of the critical span) Beaches constantly full of locals and tourists. C) (incorrect, related to a different span) 3,000 luxury hotels. D) (no support in the passage) Highly-regarded restaurants.
True Question: What is among the primary attractions of Lemnos?
Model New Participant Generated Questions New Text Generated Questions DalEye-Llama (A) What is among the primary attractions of Lemnos? (A) What is one thing to note about the eponymous village? (B) What will happen in spring 2013? (B) Who is quoted as saying that hotels benefit from the increase in the number of tourists in the area? (C) Why does Lemnos have to pay the first round of tax increases? (C) What does the quote convey about the current economic situation in Greece? (D) What is one reason why the wealthy tend to spend more time on the islands? (D) What is true of the other Greek islands? Gemini Zero-Shot (A) What can you do on the wild beaches of Lemnos? (A) What does Lemnos have? (B) Where can you swim and sunbathe almost alone? (B) What kind of nightlife does Lemnos have? (C) Why is Lemnos in line for the first round of tax increases? (C) When will Lemnos have to pay the first round of tax increases? (D) (no example) (D) Are the beaches on Lemnos crowded in the summer? Gemini Few-Shot (A) What is among the primary attractions of Lemnos? (A) What does Lemnos have dozens of? (B) (no example) (B) (no example) (C) (no example) (C) Why is Lemnos in line for the first round of tax increases? (D) Why does Lemnos have to pay the first round of tax increases? (D) Why will Lemnos have to pay the first round of tax increases? DalEye-GPT (A) What is among the primary attractions of Lemnos? (A) What does Lemnos have dozens of? (B) (no example) (B) What is true of Rhodes? (C) Why does Lemnos have to pay the first round of tax increases? (C) What does Dimitris Marinakis say about the current economic situation on Lemnos? (D) Why does Lemnos have to pay the first round of tax increases? (D) What is one thing Lemnos’ councilors say they will do?
G.2.3 DalEye-Llama: Prompt and Eye Movement Representation Variations
  1. 1.
    Fixations: Fixation Duration + Next saccade direction
    • Format: a list of fixation-level features: [fixated word index, fixated word, fixation duration in ms, direction of next saccade (backward to earlier word / within word / forward to later word)]
    • Example: [[4, ‘‘fox’’, 220, backward], ...]
  2. 2.
    • Example: [[4, ‘‘fox’’, 320, 2, 1, 3, 0], ...]
  3. 3.
    Words+Fixations
We further experiment with an alternative prompt that de-emphasizes the experimental setup and places more focus on the generation task itself:
Task Description:
Input Format:
You will receive the paragraph and eye movement data formatted as {FORMAT}
Expected Output:
Instructions:
Produce only the exact original question as your output.
{PARAGRAPH}
{SCANPATH}
@@ G.2.3 DalEye-Llama: Prompt and Eye Movement Representation Variations
In Table˜17 we present the question selection results for DalEye-Llama across the two proposed task prompts and the three input representations. Across all three task variations, Main task prompt with fixation-level input yields the best accuracy. Both Alternative task prompt and the word-level/combined inputs underperform this setting, where the choice of representation matters more than the task prompt. In Table˜18 we present goal reconstruction results that similarly suggest that the alternative prompt and eye movements representations for DalEye-Llama weaken its performance in most cases.
@@
Task Prompt Representation All Spans Diff Span Same Span
@@
Main task prompt Fixation-level $\textbf{39.2}\pm\textbf{0.9}^{***}$ $\textbf{61.4}\pm\textbf{0.9}^{***}$ $\textbf{52.5}\pm\textbf{0.4}^{***}$
@@
Alternative task prompt Fixation-level $37.2\pm 0.6^{***}_{\text{+++}}$ $60.3\pm 0.6^{***}$ $51.0\pm 0.3_{\text{+}}$
@@
Table 18: Test evaluations of the DalEye-Llama model with the fixation-level, the word-level, and combined word- and fixation-level input representations, on two task prompt variants.
New Participant New Item New Participant & Item
Question Word accuracy
Main task prompt Fixation-level $\mathbf{0.674\pm 0.013}$ $\mathbf{0.545\pm 0.017}$ $\mathbf{0.548\pm 0.023}$
Word-level $0.185\pm 0.019$ $0.128\pm 0.014$ $0.125\pm 0.016$
Both fixation- and word-level $0.191\pm 0.020$ $0.130\pm 0.014$ $0.134\pm 0.016$
Alternative task prompt Fixation-level $0.635\pm 0.014$ $0.534\pm 0.017$ $0.547\pm 0.023$
Word-level $0.182\pm 0.018$ $0.129\pm 0.014$ $0.133\pm 0.016$
@@
Question Category accuracy
Main task prompt Fixation-level $\mathbf{0.558\pm 0.014}$ $0.450\pm 0.017$ $0.428\pm 0.023$
Word-level $0.448\pm 0.015$ $0.432\pm 0.015$ $0.438\pm 0.022$
Both fixation- and word-level $0.286\pm 0.015$ $0.257\pm 0.015$ $0.259\pm 0.020$
Alternative task prompt Fixation-level $0.550\pm 0.015$ $0.484\pm 0.016$ $0.473\pm 0.024$
Word-level $0.537\pm 0.018$ $\mathbf{0.538\pm 0.019}$ $\mathbf{0.529\pm 0.024}$
Both fixation- and word-level $0.515\pm 0.018$ $0.507\pm 0.018$ $0.498\pm 0.024$
@@
Main task prompt Fixation-level $\mathbf{0.556\pm 0.008}$ $0.253\pm 0.005$ $0.244\pm 0.006$
Word-level $140\pm 0.014 0.0$$89\pm 0.005 0.0$$085\pm 0.006$
Both fixation- and word-level $146\pm 0.015 0.0$$88\pm 0.005 0.0$$085\pm 0.006$
Alternative task prompt Fixation-level $0.493\pm 0.009$ $\mathbf{0.264\pm 0.004}$ $\mathbf{0.262\pm 0.007}$
Word-level $126\pm 0.013 0.0$$77\pm 0.005 0.0$$073\pm 0.006$
Both fixation- and word-level $096\pm 0.012 0.0$$65\pm 0.004 0.0$$063\pm 0.005$
@@
Main task prompt Fixation-level $\mathbf{0.776\pm 0.004}$ $0.631\pm 0.003$ $0.626\pm 0.004$
Word-level $0.429\pm 0.012$ $0.410\pm 0.007$ $0.408\pm 0.009$
Both fixation- and word-level $0.426\pm 0.013$ $0.406\pm 0.007$ $0.402\pm 0.010$
Alternative task prompt Fixation-level $0.746\pm 0.005$ $\mathbf{0.633\pm 0.003}$ $\mathbf{0.632\pm 0.004}$
Word-level $0.410\pm 0.013$ $0.388\pm 0.007$ $0.384\pm 0.009$
Both fixation- and word-level $0.409\pm 0.010$ $0.393\pm 0.005$ $0.390\pm 0.006$